# FairJob: A Real-World Dataset for Fairness in Online Systems

### v1

@@ A.1 Data collection and use
As illustrated in Figure 6, the process starts with users navigating Publisher and Advertiser websites (typically newspapers and retailer shops respectively). Upon user consent [^10] In accordance with GDPR and Article 82 of the Data Protection Act that require the provider to ask consent of data subjects if it was reading/writing information to the user’s device., user information about the events such as visits or product views [^11] The original data does not contain any ”special categories” of personal data listed under Article 9 of the GDPR, processing of which is prohibited, except in limited circumstances set out in Article 9 of the GDPR. are collected and identified by means of browser cookies. Users are subject to personalized advertising (if the job campaign was chosen and the display opportunity was won) until the end of the data collection period. Subsequently, only won displays coming from Publisher and Advertiser partners are joined by cookie identifier on the AdTech platform to form the raw dataset, dropping cookie ids when they are not needed anymore. To ensure confidentiality, the data has been sub-sampled non-uniformly to avoid disclosing business metrics. Feature names have been anonymized, and their values randomly projected to preserve predictive power while rendering the recovery of the original features or user context practically impossible [^12] The orginal data is still used in the AdTech company, however, re-indentification of the pseudomymized data is impossible due to additional randomization techniques during data anonymization to comply with Article 4(5) of GDPR.. The dataset does not contain the relevant attributes such as gender; however, it includes a gender proxy which we discuss in detail in the main text.
@@
The data is released under the CC-BY-NC-SA 4.0 license which gives liberty to Share and Adapt this data provided that the respect of the Attribution, NonCommercial and ShareAlike conditions. We focus in this paper on algorithmic fairness analysis as a specific use-case, but it does not have to restricted to this. For example, the dataset can be used as a baseline for improving deep learning models on tabular data and to study methods on embedding creation for numerical and categorical features as done by Gorishniy et al. [2021, 2022], Grinsztajn et al. [2022], Shwartz-Ziv and Armon [2022], Matteucci et al. [2023]. Another possible usage is to explore privacy-preserving techniques for tabular data as in Donhauser et al. [2024]. Additionally, this dataset can be used to improve the generation techniques on tabular data, as described in [Jesus et al., 2022, van Breugel and van der Schaar, 2024]. We compare FairJob dataset details to other commonly used open-source tabular datasets further in Section D. And lastly, as we discussed in the main text, FairJob datasets presents a challenge of training fair and robust models on strongly imbalanced data [Jesus et al., 2022, Yang et al., 2024].
A.2 Dataset detailed description
@@
  • impression_id is a unique identifier assigned to each impression, i.e. online session that can have several products at the same time.
  • cat0, $\dots$, cat5 are anonymized categorical user features.
  • cat6, $\dots$, cat12 are anonymized categorical product features.
@@
  • senior is a binary feature that describes the seniority of the job position, i.e. an assistant role is 0, a managerial role is 1. This feature was created during data processing step from the product title feature: if the product title contains words describing managerial role (e.g. ’president’, ’ceo’, and others), it is assigned to 1, otherwise to 0.
@@
In Figures 9 and 10 we plot correlation matrix between numerical features in the original dataset and open-sourced dataset FairJob, where we added noise for anonymization. We observe that correlations are generally slightly weaker for most features, when observing the anonymized data, that can be better seen in Figure 11, where we plot difference between the correlations. This fact comes naturally from the added noise, however, lower correlation values might translate into higher classification difficulty.
@@
A.3 Dataset statistics
Table 1, 2 and 3 demonstrate the available statistics for categorical features cardinalities, index features and binary features respectively. Additionally, we provide statistics for selection bias of the job campaign that comes from the Criteo AdTech company (outside of FairJob dataset) in Table 4, which we take into account further to compute utility metrics.
@@ B.1 Metrics
Demographic parity requires equal proportion of positive predictions in each group ("No Disparate Impact") which in our case can be translated as same proportion of shown senior job ads in each group of the protected attribute. Demographic parity can be thought of as a stronger version of the US Equal Employment Opportunity Commission’s "four-fifths rule", which requires that the “selection rate for any race, sex, or ethnic group [must be at least] four-fifths (4/5) (or eighty percent) of the rate for the group with the highest rate [^13] See the Uniform Guidelines on Employment Selection Procedures, 29 C.F.R. §1607.4(D) (2015)..
Equalized odds [Zafar et al., 2017] ensures that a machine learning model works equally well for different groups. In case of job ads, equalized odds would force the prediction of both positive and negative outcomes to be the same which might generate more false positive predictions for one group versus others, resulting in worse outcome. Equal of opportunity [Hardt et al., 2016] ensures that a machine learning model works equally well only on positive outcomes for different groups which results in capturing the costs of missclassification disparities.
@@
The loss function is log-loss and report it as NLLH (negative log-likelihood). We also report AUC (Area under the ROC Curve) as a description of prediction power on strongly imbalanced data in binary classification problems. Additionally, we consider click-rank utility $U$ and product-rank utility for biased data $\tilde{U}$, proposed in the main text Section 3.4.
@@
Fairness methods.
To study possible trade-off between utility and fairness, we use in-processing fairness methods of adding fairness-inducing penalty to a loss function during training [Kamishima et al., 2011]:
$\hat{Y}=\arg\min\ \ \mathcal{L}(\hat{Y},Y)+\lambda\cdot\text{Penalty}(\hat{Y},Y,A).$ (7)
Parameter $\lambda$, or fairness_multiplier, controls the trade-off between the model’s predictive accuracy $\mathcal{L}(\hat{Y},Y)$ and fairness. Adjustments on $\lambda$ allows to control the importance of fairness relative to accuracy. These methods remove the influence of protected attribute on the model’s output without restrictions on the data [Kamishima et al., 2011, Bechavod and Ligett, 2017, Mary et al., 2019]. We implement a penalty based on the approach described in Bechavod and Ligett [2017].
B.3 Results
Baseline models.
We train (i) a Dummy in the unaware regime to obtain the first baseline and (ii) XGB in two regimes (unaware and unfair) to achieve a more reasonable baseline performance, see Table 8. We notice that the Dummy classifier is perfectly fair with respect to DP which is reasonable since it did not learn at all as can be seen from the NLLH results. However, the utility metrics $U$ and $\tilde{U}$ of Dummy do not differ much from XGB which is due to very strong imbalance in the data. We remark that $\tilde{U}$ is expectedly higher for fairer (unaware) models, while $U$ is better for the unfair model. There is a slight difference in terms of NLLH and AUC-ROC for XGB models, and higher $U$ corresponds to higher AUC-ROC.
NLLH $\downarrow$ AUC $\uparrow$ DP $\downarrow$ $U$uparrow $\tilde{U}$uparrow
Dummy unaware 0.69239 0.50000 0.00000 0.01009 0.01245
XGB unaware 0.05491 0.75787 0.00278 0.01017 0.01276
XGB unfair 0.05736 0.76201 0.00323 0.01037 0.01236
The XGB models were used for feature importance measurement reported in Figure 7. We observed that the results reported in Figure 7 for both unfair and unaware cases were stable when changing the seed.
Fairness-utility trade-off.
B.4 Possible improvements.
@@ Appendix D Comparison to other tabular datasets
COMPAS 7214 14 37 binclass sex, race
Gesture Phase 9873 32 0 multiclass -
Churn Modelling 10000 10 1 binclass -
California Housing 20640 8 0 regression income
House 16H 22784 16 0 regression -
Adult 48842 6 8 binclass sex, race, gender
Otto Group Products 61878 93 0 multiclass -
Higgs Small 98049 28 0 binclass -
Diabetes 101766 10 40 binclass gender
Facebook Comments Volume 197080 50 1 regression -
Santander Customer Transactions 200000 200 0 binclass -
KDD Cencus-Income 299285 34 7 binclass sex, race
Covertype 581012 54 0 multiclass -
[0pt][0pt] FairJob 1072226 36 18 binclass gender
MSLR-WEB10K (Fold 1) 1200192 136 0 regression -

### v2

@@ A.1 Data collection and use
As illustrated in Figure 7, the process starts with users navigating Publisher and Advertiser websites (typically newspapers and retailer shops respectively). Upon user consent [^10] In accordance with GDPR and Article 82 of the Data Protection Act that require the provider to ask consent of data subjects if it was reading/writing information to the user’s device., user information about the events such as visits or product views [^11] The original data does not contain any ”special categories” of personal data listed under Article 9 of the GDPR, processing of which is prohibited, except in limited circumstances set out in Article 9 of the GDPR. are collected and identified by means of browser cookies. Users are subject to personalized advertising (if the job campaign was chosen and the display opportunity was won) until the end of the data collection period. Subsequently, only won displays coming from Publisher and Advertiser partners are joined by cookie identifier on the AdTech platform to form the raw dataset, dropping cookie ids when they are not needed anymore.
Further, the data is collected by trackers upon a page call and is commonly stored as a tuple of page and client information, which we call a click. Several steps are in place to ensure that clicks on ads must result from a human user with genuine interest. This steps are stated in the guidelines that are accepted by publishers [^12] The guidelines are available on the official webpage: https://www.criteo.com/supply-partner-guidelines/.. Any method or mechanism that artificially generates clicks or impressions is strictly prohibited, including but not limited to the following:
  • Clicks generated by publishers clicking on their own ads.
  • Repeated clicks on the same ad unit.
  • Publishers encouraging users to click on their ads (examples may include: any language encouraging users to click on ads; ad implementations that may cause a high volume of accidental clicks; monetary or non-monetary incentives or rewards for clicks, etc.).
  • Use of automated clicking tools or traffic sources, robots, or other deceptive software, click spam or click injections.
  • Use of any other artificial mechanism to generate or inflate clicks.
  • Clicks generated outside of the ad surface will not be counted as intended clicks.
Additional safeguards are related to clearly stating that the user sees an ad. Any native ad formats displayed on the publisher’s site must be clearly marked as an ad. Ads should not be placed very close to or underneath buttons or any other object such that the placement of the ad interferes with a user’s typical interaction with the Site content or functionalities.
Publishers accept the guidelines which state the following:
  • Publishers may offer their users the opportunity to view ads in exchange for user rewards or incentives, but only if the user is not forced or incentivized to interact with the ad, such as incentivized click.
  • Publishers must not, directly or indirectly, provide incentives to users in exchange for clicks on ads, or make use of any mechanism or monetizable reward to incentivize clicks.
To ensure confidentiality, the collected data has been sub-sampled non-uniformly to avoid disclosing business metrics. Feature names have been anonymized, and their values randomly projected to preserve predictive power while rendering the recovery of the original features or user context practically impossible [^13] The orginal data is still used in the AdTech company, however, re-indentification of the pseudomymized data is impossible due to additional randomization techniques during data anonymization to comply with Article 4(5) of GDPR.. The dataset does not contain the relevant attributes such as gender; however, it includes a gender proxy which we discuss in detail in the main text.
@@
The data is released under the CC-BY-NC-SA 4.0 license which gives liberty to Share and Adapt this data provided that the respect of the Attribution, NonCommercial and ShareAlike conditions. We focus in this paper on algorithmic fairness analysis as a specific use-case in job-seeking contexts, but it does not have to restricted to this. Fairness in job advertising is a crucial step towards ensuring fairness in any advertising campaign involving sensitive topics. The FairJob datasets presents a challenge of training fair and robust models on strongly imbalanced data [Jesus et al., 2022, Yang et al., 2024]. While many of the features in the FairJob dataset (such as user features and categorical product features) are not specific to job advertising and can be applied to other domains for improving click-prediction methods, however, the consumer behavior may vary significantly from a job seeker’s behavior. Additionally, the dataset can be used as a baseline for improving deep learning models on tabular data and to study methods on embedding creation for numerical and categorical features as done by Gorishniy et al. [2021, 2022], Grinsztajn et al. [2022], Shwartz-Ziv and Armon [2022], Matteucci et al. [2023]. Another possible usage is to explore privacy-preserving techniques for tabular data as in Donhauser et al. [2024]. Moreover, this dataset can be used to improve the generation techniques on tabular data, as described in Jesus et al. [2022], van Breugel and van der Schaar [2024]. We compare FairJob dataset details to other commonly used open-source tabular datasets further in Section D.
A.2 User privacy protection
The original dataset does not include any data that directly identifies a user, such as names, postal addresses, or email addresses in plain text, so the original data is pseudonymized. However, due to possible uniqueness in high-dimensional datasets, it has been shown that some records in pseudonymized datasets might be re-identifiable by an adversary. We describe a possible re-identification process by an adversary and demonstrate that it is practically impossible.
In the FairJob dataset, there are both user and product features. We assume that only user features could potentially be used to identify a user. The product features are categorical, based on internal company categories across different products and brands, or on information provided by clients, and thus are not related to users.
We can suppose that an adversary runs a company that collects user browsing features and, therefore, has an auxiliary dataset with user information. The adversary’s goal is to identify a user from the FairJob dataset. Since all features are anonymized, the adversary cannot directly match the features and must first align FairJob feature distributions with the adversary’s company’s feature distributions and then find a unique association from a FairJob record to the adversary’s company record, thereby identifying the user. However, since we anonymized, added noise, and standardized continuous user features, re-identification becomes impossible as the original scale is not provided. The only features of potential interest to the adversary are two categorical user features.
The categorical user features are based on internal company categories, which are unlikely to be known to the adversary due to business confidentiality, making it impossible for the adversary to reconstruct them. Additionally, features cat0 and cat1 have the same cardinality, and their distributions are not significantly different, making them indistinguishable and effectively reducing the number of possible unique variations. Most importantly, both features cat0 and cat1 have a cardinality of 9, meaning the tuples (cat0, cat1) can have a maximum of 81 variations among 1 million records. In FairJob, each present variation of (cat0, cat1) corresponds to at least 10 records, making it impossible for an instance (cat0, cat1) to serve as a pseudonym for a user, i.e., user re-identification based on these two features is not possible.
All these measures make it impossible to re-identify users from the FairJob dataset.
A.3 Dataset detailed description
@@
  • impression_id is a unique identifier assigned to each impression, i.e. online session that can have several products at the same time. An impression, or successful online session, is recorded when a product banner (or multiple banners) is loaded and begins to render on the publisher’s page. An impression can subsequently lead to a click event.
  • cat0, $\dots$, cat1 are anonymized categorical user features.
  • cat2, $\dots$, cat12 are anonymized categorical product features.
@@
  • senior is a binary feature that describes the seniority of the job position, i.e. an assistant role is 0, a managerial role is 1. It includes both managerial and individual contributor (IC) roles, as our aim was to encompass any senior-level positions. This feature was created during the data processing step from the product title feature: if the product title contains words describing managerial role (e.g. ’president’, ’ceo’, and others), it is assigned to 1, otherwise to 0.
@@
In Figures 10 and 11 we plot correlation matrix between numerical features in the original dataset and open-sourced dataset FairJob, where we added noise for anonymization. We observe that correlations are generally slightly weaker for most features, when observing the anonymized data, that can be better seen in Figure 12, where we plot difference between the correlations. This fact comes naturally from the added noise, however, lower correlation values might translate into higher classification difficulty.
@@
A.4 Dataset statistics
Table 2, 3 and 4 demonstrate the available statistics for categorical features cardinalities, index features and binary features respectively. Additionally, we provide statistics for selection bias of the job campaign that comes from the Criteo AdTech company (outside of FairJob dataset) in Table 5, which we take into account further to compute utility metrics.
@@ B.1 Metrics
Demographic parity requires equal proportion of positive predictions in each group ("No Disparate Impact") which in our case can be translated as same proportion of shown senior job ads in each group of the protected attribute. Demographic parity can be thought of as a stronger version of the US Equal Employment Opportunity Commission’s “four-fifths rule”, which requires that the “selection rate for any race, sex, or ethnic group [must be at least] four-fifths (4/5) (or eighty percent) of the rate for the group with the highest rate” [^14] See the Uniform Guidelines on Employment Selection Procedures, 29 C.F.R. §1607.4(D) (2015)..
Equalized odds [Zafar et al., 2017] ensures that a machine learning model works equally well for different groups. In case of job ads, equalized odds would force the prediction of both positive and negative outcomes to be the same which might generate more false positive predictions for one group versus others, resulting in worse outcome. Equal of opportunity [Hardt et al., 2016] ensures that a machine learning model works equally well only on positive outcomes for different groups which results in capturing the costs of misclassification disparities.
@@
The loss function is log-loss and report it as NLLH (negative log-likelihood). We also report AUC (Area under the ROC Curve) as a description of prediction power on strongly imbalanced data in binary classification problems. Additionally, we consider click rank utility $U$ and product-rank utility for biased data $\tilde{U}$, proposed in the main text Section 3.4.
@@
B.3 Possible improvements.
@@ Appendix D Comparison to other tabular datasets
D.1 Click-prediction datasets
We argue that the FairJob dataset is representative enough to evaluate click prediction algorithms, based on three main points:
  • real-world data – the dataset is based on real-world data and contains all the necessary information. We have provided detailed documentation to support this.
  • comparative size – the dataset is comparable in size to other widely-used specialized recommendation datasets.
  • baseline experiments – we performed baseline experiments that produced reasonable results with respect to both performance and the utility-fairness trade-off.
The dataset contains pseudonymized users’ context and publisher features collected from a job-targeting campaign run over five months. Therefore, the dataset contains slices of real-world data.
To properly compare the FairJob dataset to other click prediction datasets, we searched for the most used datasets for click prediction. We identified the following widely-used datasets for click prediction based on benchmark papers [Zhu et al., 2021, Mao et al., 2023]:
  • Criteo 2014 Dataset. It consists of ad click data over a week and comprises 26 categorical feature fields and 13 numerical feature fields.
  • Avazu 2015 Dataset. It contains 10 days of click logs and has a total of 23 fields with categorical features including app id, app category, device id, etc.
Both datasets are sampled from real click logs in production and contain tens of millions of samples. However, both lack explicit user and item field information. In addition to click prediction datasets, other recommendation system problems of similar complexity, such as MovieLens for predicting engagements and Frappe related to app usage, have datasets comparable in size to FairJob [Zhu et al., 2021, Mao et al., 2023]. See the detailed comparison in Table 9.
Table 9: Comparison of FairJob to other most frequently used tabular datasets for click prediction and recommendation system problems of similar complexity.
dataset  # rows  # features
Criteo 45,840,617 39
Avazu 40,428,967 22
MovieLens 2,006,859 3
[0pt][0pt] FairJob 1,072,226 60
Frappe 288,609 10
We acknowledge that the FairJob dataset does not contain all features typically used for training click prediction models in the industry, due to business confidentiality. However, despite hand-picking the most relevant features, our dataset still contains 60 features, which is the largest number of real features in a dataset for click prediction.
The FairJob dataset does focus on a subproblem of click prediction within the context of job advertising campaigns, where users are pre-selected as having potential job-seeking profiles. Consequently, the dataset size is represented by a smaller number of records specific to this campaign. The click rate of 0.007% is consistent with the click rates observed in larger-scale datasets.
Furthermore, 1 million records for an unbalanced classification problem seems a reasonable size compared to other tabular datasets, as also stated in the next subsection. Our experiments in Appendix B demonstrate that baseline classification algorithms can successfully predict clicks on the test set.
D.2 Fairness-aware datasets
@@
COMPAS 7,214 14 37 binclass sex, race
Gesture Phase 9,873 32 0 multiclass -
Churn Modelling 10,000 10 1 binclass -
California Housing 20,640 8 0 regression income
House 16H 22,784 16 0 regression -
Adult 48,842 6 8 binclass sex, race, gender
Otto Group Products 61,878 93 0 multiclass -
Higgs Small 98,049 28 0 binclass -
Diabetes 101,766 10 40 binclass gender
Facebook Comments Volume 197,080 50 1 regression -
Santander Customer Transactions 200,000 200 0 binclass -
KDD Cencus-Income 299,285 34 7 binclass sex, race
Covertype 581,012 54 0 multiclass -
[0pt][0pt] FairJob 1,072,226 36 18 binclass gender
MSLR-WEB10K (Fold 1) 1,200,192 136 0 regression -
@@ Appendix E Broader impact
Click prediction is a fundamental task in online advertising and recommendation systems, as noted in recent benchmarking papers [Zhu et al., 2021, Mao et al., 2023]. This is an active research field where improvements in click prediction algorithms lead to improvements in recommendation systems in general [Mao et al., 2023]. However, improvements in general models do not always translate to specialized models, thus, there is a need to verify algorithms directly with available resources. The FairJob dataset represents a particular case of the click prediction problem, focusing on job-seeking user profiles and job advertising product features. However, many features (user features and categorical product features) are not campaign-specific and can be found in other applications outside job advertising. By open-sourcing the FairJob dataset, we provide access to a real-world problem of finding a trade-off between utility and fairness in job offer advertising. We argue that advances in fair machine learning methods should be validated in real-world scenarios, as the distribution of features and their impact on outputs and fairness can vary significantly across different applications.
As described in Section 3, advertising, like almost any complex problem, is subject to various biases [Hort et al., 2023]. We have clearly stated the potential bias sources, analyzed their impact, and discussed possible corrections. While we cannot definitively prove that the properties of the FairJob dataset can be extended to every click prediction problem, it serves as a representation of a specific real-world scenario where the fairness-utility trade-off is encountered. If the FairJob dataset helps to find better methods for improving fairness in job advertising, these methods might also be applicable to other advertising campaigns, as similar user signals may be present in the features used to predict clicks or to associate with the utility-fairness trade-off.